Someone please answer this question immediately..
A cube of ice of edge 4cm is placed in an empty cylindrical glass of inner diameter 6 cm. Assume that ice melts uniformly from each side so that it always retains cubical shape. Remember that ice is lighter than water. Find the length of the edge of the ice cube at the instant it just leaves contact with the bottom of the glass.

Grade:11

1 Answers

Arun
5 years ago
Sol. Let x → edge of ice block When it just leaves contact with the bottom of the glass. h→ height of water melted from ice W = U ⇒x3 * pice * g = x2 * h * pw * g Again, volume of water formed, from melting of ice is given by, 43 – x3 = π * r2 * h – x2h ( because amount of water = (πr2 – x2)h) ⇒43 – x3 = π * 32 * h – x2h Putting h = 0.9 x ⇒x = 2.26 cm  RegardsArun